OpenStudy (anonymous):

help me! please log_{5}5^{8x+1}

OpenStudy (campbell_st):

so what do you need to do to the problem

OpenStudy (anonymous):

simplify

OpenStudy (campbell_st):

ok... so when the base of the log and base of the exponent are the same you just get the power \[\log_{a} a^x = a\] the log is base a, and the exponential is base a, the the answer is just the power. Same applys in your question log base is 5 and the base of the exponential is 5 hope this helps

OpenStudy (anonymous):

mm no this no helped me but thanx

OpenStudy (campbell_st):

do you have a calculator...?

OpenStudy (anonymous):

yes

OpenStudy (campbell_st):

ok.... whats the value of \[\log(10^5)\] and whats the value of \[\ln(e^4) = \]

OpenStudy (anonymous):

5 and 4?

OpenStudy (campbell_st):

ok so look at the problem you jsut solved